OpenStudy (anonymous):

Show how to solve seven-ninths times five-fourteenths with and without cancellation.

OpenStudy (anonymous):



OpenStudy (anonymous):

now simplify

OpenStudy (anonymous):

7/9 * 5/14 = (7*5) / (9*14) = 35/126 which reduces to 5/18 -- without cancelling Let me draw the one for cancelling....